OpenStudy (anonymous):

The leg of a right triangle is 4 units and the hypotenuse is 9 units. What is the length, in units, of the other leg of the triangle?

OpenStudy (anonymous):

Hi Welcome to Openstudy! Do you know the Pythagorean theorem? \(\sf \large hypotenuse^2=adjacent^2+opposite^2\) or \(\sf \large c^2=a^2+b^2\) you are looking for the other leg(assume it is leg b), so you just have to transfer a to be with c so you have to subtract both sides by \(\sf a^2\) and you'll get \(\sf \large b^2=c^2-a^2\) where c = 9 units and a= 4 units plug in your given values and solve for b (:

OpenStudy (anonymous):

5

OpenStudy (anonymous):

THank you so much :)

OpenStudy (anonymous):

both of you